OpenStudy (anonymous):

1.Write the equation of the line which passes through (5, –2) and is parallel to x = 4. (1 point)

OpenStudy (ash2326):

The line x=4 has a slope of infinity so line parallel to it'll also have slope of infinity Standard equation of a line with slope m and passing throgh (x1,y1) (y-y1)/(x-x1)=m y-y1=m(x-x1) here m= infinity or let m= 1/p where p=0 so (x1,y1)=(5, -2) \[y+2=\frac{1}{p}(x-5)\] multiplying p to both sides we get \[p(y+2)=x-5\] p=0 so \[0=x-5\] so the line is \[x=5\]

OpenStudy (ash2326):

Sassyjazz did you understand this?

OpenStudy (anonymous):

yes i understand